OpenStudy (anonymous):

5(x - 3) + 4(x + 3) = 3(x - 1)

OpenStudy (anonymous):

5(x - 3) + 4(x + 3) = 3(x - 1) 5x - 15 + 4x + 12 = 3x - 3